OpenStudy (anonymous):

a 100 kg person stands on a scale in an elevator accelerating downwards at 2 m/s/s. what will the scale read as their apparent weight

OpenStudy (michele_laino):

we can see the elevator like an inertial system wherein is present a fictitious force like in the subsequent drawing:  so the apparent weight \(W\) is: \(W=mg-ma=m(g-a)=...?\)